# Granular-Ball-Induced Multiple Kernel K-Means

### v1

@@ 1 Introduction
  • The proposed GBK effectively reduces storage requirements and computational costs while ensuring superior clustering performance because it allows for reducing the kernel matrix size from “the number of samples $n$” to “the number of granular-balls $m$” ($m\ll n$).
@@ 3.3 Multiple Kernel Fusion
Input: Granular-ball kernel matrix $\{\bar{\mathbf{K}}_{k}^{GB}\}_{k=1}^{\omega}$, initial ${t}=1$.
@@
2:  while $flag$ holds do
@@ 3.4 GB-induced Multiple Kernel K-means
To balance the influence of different kernels and ensure the rationality of sample assignments, we optimize the model by minimizing the following objective for the kernel coefficient $\sigma_{k}$ ($k=1,2,...,w$):
@@ 4 Experiments
In this section, we conduct the following experiments, including evaluating the clustering performance of the proposed granular-ball-induced multiple kernel clustering against recent strong multiple kernel baselines on nine single-view datasets and three multiple-view datasets; comparing different anchor-sampling methods in multi-kernel clustering frameworks to verify the effectiveness of using granular-ball computing; and analyzing optimization convergence and computational cost. The experiments are conducted using MATLAB R2022a  [^1] Source code and supplementary materials are available at: https://github.com/WangYifan4211115/GB-MKKM..
@@ Acknowledgments
This research was supported in part by the National Natural Science Foundation of China under Grant Nos. 62221005, 62450043, 62222601, and 62176033.

### v2

@@ 1 Introduction
  • The proposed GBK effectively reduces storage requirements and computational costs while ensuring superior clustering performance because it allows for reducing the kernel matrix size from “the number of samples nn” to “the number of granular-balls mm” ($m\ll n$).
@@ 3.3 Multiple Kernel Fusion
Input: Granular-ball kernel matrix $\{\bar{\mathbf{K}}_{k}^{GB}\}_{k=1}^{\omega}$, initial t=1{t}=1.
@@
2:  while flagflag holds do
@@ 3.4 GB-induced Multiple Kernel K-means
To balance the influence of different kernels and ensure the rationality of sample assignments, we optimize the model by minimizing the following objective for the kernel coefficient $\sigma_{k}$ (k=1,2,,wk=1,2,...,w):
@@ 4 Experiments
In this section, we conduct the following experiments, including evaluating the clustering performance of the proposed granular-ball-induced multiple kernel clustering against recent strong multiple kernel baselines on nine single-view datasets and three multiple-view datasets; comparing different anchor-sampling methods in multi-kernel clustering frameworks to verify the effectiveness of using granular-ball computing; and analyzing optimization convergence and computational cost. The experiments are conducted using MATLAB R2022a  [^1] Source code and supplementary materials are available at: https://github.com/WangYifan4211115/GB-MKKM.
@@ Acknowledgments
This research was supported by the Chongqing Graduate Research Innovation Program CYB25259 and the National Natural Science Foundation of China under Grant Nos. 62221005, 62450043, 62222601, and 62176033.